### v1

# A Semantic Search Engine for Mathlib4

Abstract
The interactive theorem prover, Lean, enables the verification of formal mathematical proofs and is backed by an expanding community. Central to this ecosystem is its mathematical library, mathlib4, which lays the groundwork for the formalization of an expanding range of mathematical theories. However, searching for theorems in mathlib4 can be challenging. To successfully search in mathlib4, users often need to be familiar with its naming conventions or documentation strings. Therefore, creating a semantic search engine that can be used easily by individuals with varying familiarity with mathlib4 is very important. In this paper, we present a semantic search engine333Our search engine is expected to launch within a month, available as a cloud service. for mathlib4 that accepts informal queries and finds the relevant theorems. We also establish a benchmark for assessing the performance of various search engines for mathlib4.
keywords:
Semantic Search, Machine Learning, Benchmarkscategory:
[…]
Lean [2, 3] is an interactive theorem prover built on dependent type theory, designed to verify mathematical proofs written in a formal language and thereby enhancing their rigor. It has a vibrant and supportive community, with its popularity growing among mathematicians. A prime example of the Lean community’s collaborative spirit is mathlib4, the mathematical library for Lean 4. This library, regularly updated by contributors from around the globe, acts as a basis for the formalization of new mathematical theories. This eliminates the need to repeatedly formalize established results, as users can simply check if mathlib4 contains the necessary theorems referenced in informal proofs. However, locating these theorems is often challenging due to the limitations of officially provided search tools, which struggle to find relevant theorems with informal queries.
There are primarily two methods to search for theorems in mathlib4: using the mathlib4 documentation444https://leanprover-community.github.io/mathlib4_docs/ and searching the GitHub repository of mathlib4555https://github.com/leanprover-community/mathlib4. The mathlib4 documentation allows users to search for theorems by their formal names. However, this feature can be challenging for beginners to utilize effectively, as they may not be familiar with the naming conventions. For example, Cauchy’s Mean Value Theorem is named as exists_ratio_deriv_eq_ratio_slope in mathlib4, meaning a direct search for "Cauchy’s Mean Value Theorem" yields no results. An alternative method involves searching within the GitHub repository of mathlib4, which allows for keyword-based searches across the source files, including formal statements, proofs, and documentation strings. This method can locate Cauchy’s Mean Value Theorem as it is mentioned in the documentation string of exists_ratio_deriv_eq_ratio_slope. However, this approach faces two issues: many theorems in mathlib4 lack documentation strings, and semantically similar user queries that don’t exactly match the theorems or documentation strings may lead to unsuccessful searches.
Consequently, neither method adequately supports finding theorems based on informal queries, leading to beginners spending significant time on this task. According to a survey among students working on formalizing convex analysis at Peking University, they spend half of their time searching for theorems. Moreover, conversations on Zulip666https://leanprover.zulipchat.com/#narrow/stream/219941-Machine-Learning-for-Theorem-Proving/topic/Semantic.20Search.20for.20Mathematics have highlighted the need for creating a semantic search engine for mathlib4. Consequently, the development of such a search engine for mathlib4 is highly desirable to enhance the efficiency of theorem retrieval.
In this paper, we introduce a semantic search engine for mathlib4 that allows users to input an informal query and retrieve a list of relevant theorems from mathlib4. To construct this search engine, we translate the formal statements of mathlib4 theorems into informal ones and integrate these pairs into a database. Each database entry consists of a formal theorem statement and its informal version. Upon receiving a user query, we augment the query for improved context understanding and perform semantic search across the database to find relevant results. Additionally, we have established a benchmark to compare the effectiveness of various search engines for mathlib4.
The remaining part of the paper is organized as follows. In Section 2, we review the related works on text retrieval and mathematical information retrieval. Section 3 describes our approach to developing a semantic search engine for mathlib4. Section 4 presents the mathlib4 semantic search benchmark. Numerical results are discussed in Section 5. We conclude this paper in Section 6.
2 Related Work
Text Retrieval.
Text retrieval is the task of finding relevant information within a corpus based on user queries. Early methods, such as BM25 [26, 25], used sparse vector representations for queries and documents, assessing relevance by comparing these vectors with certain weighting techniques [27]. While effective in measuring lexical similarity, these approaches fall short in capturing the semantic similarity between queries and documents.
To address this limitation, deep learning techniques [6, 33, 17, 14] have been introduced. Utilizing deep neural networks, these techniques encode queries and documents into dense vectors, thereby assessing relevance based on semantic similarity. Further developments have been made in neural architectures and training paradigms. There are two main architectures: the cross-encoder [23, 22] and the bi-encoder [9, 24, 21, 20]. Cross-encoders take the concatenation of query and document as input and produce the final relevance of this query-document pair, while bi-encoders map the query and document into vectors separately, determining relevance through similarity between the two vectors. Training paradigms have also evolved, with the Inverse Close Task (ICT) [12] initially proposed for dense retriever pre-training. Subsequently, other pre-training tasks have been developed, including Body First Selection and Wiki Link Prediction, both introduced in [1]. Recent studies [18, 30, 28, 32] have explored large-scale unsupervised pre-training using contrastive loss, followed by fine-tuning on smaller, labeled datasets. Pre-training with extensive text pairs allows the language model to grasp textual semantics and the fine-tuning stage further enhances its performance across various retrieval tasks. However, the authors of [31] argue that a two-stage training approach might not be necessary, demonstrating that directly fine-tuning a decoder-only model on both synthetic and labeled data can yield competitive results.
Mathematical Information Retrieval.
Mathematical Information Retrieval (MIR) differs from text retrieval in that it involves queries and documents that contain mathematical formulas. These formulas are highly structured, which distinguishes them from plain text. To effectively capture the semantics of math formulas, several representations are employed. The Symbol Layout Tree (SLT) [35] preserves the original layout of formulas, while the Operator Tree (OPT) [5] represents mathematical symbols as nodes, with edges denoting the relationships between operators and operands. Classical MIR methods [4, 5, 11, 36, 8, 19, 34] rely on structure search, identifying matching substructures across various features. Among these, the Approach0 structural search method [41, 38] has shown to be particularly effective. It indexes formulas through leaf-root paths in the OPT and utilizes subexpression matching to assess formula similarity. With advancements in deep learning, dense retrievers have been integrated into MIR, often combined with structural searches [8, 39, 40, 37]. A notable example is the Approach0 hybrid search [39], which combines Approach0 structure search with a bi-encoder dense retriever, ColBERT [10]. This combination not only facilitates effective formula matching but also enhances understanding of context.
The preceding literature focuses on extracting mathematical content from a corpus composed of natural language texts and formulas formatted in markup languages. This contrasts with our objective of conducting searches within mathlib4, a formal mathematical library. The work most closely related to ours is Moogle777https://www.moogle.ai/, a semantic search engine for mathlib4. However, its technical details have not been released. We will compare the performance of Moogle and our search engine in Section 5.
3 Methodology
In this section, we will describe the implementation of our semantic search engine for mathlib4. This engine is designed to accept a user query in natural language and return a list of relevant theorems in mathlib4. Our approach involves converting formal theorems from mathlib4 into their informal counterparts, as illustrated in Figure 1. These informal-formal theorem pairs are then vectorized and stored in a vector database, a step that can be executed offline. When a user query is submitted, we augment the query to better grasp its context, vectorize the enriched query, and locate the corresponding theorems in the embedding space. In the following subsections, we will elaborate on the informalization of mathlib4, the design of the search engine, and the method of query augmentation.
3.1 Informalizing Mathlib4
Our strategy for informalizing mathlib4 involves employing a large language model (LLM). Central to this strategy is providing the LLM with sufficient context to accurately grasp the formal theorem’s exact meaning. To this end, we not only extract the theorem’s name, statement, and documentation string888For many theorems in mathlib4 that lack a documentation string, we will extract only the theorem name and statement. from the mathlib4 documentation but also include related definitions found in the theorem statements through hyperlinks. For example, as illustrated in Figure 2, we extract the definition of Exists.choose because it is referenced and linked in the theorem Exists.choose_spec. The gathered information is then fed into an LLM999We use gpt-3.5-turbo-16k for generating informal names and statements, setting the temperature to . to generate informal theorem names and statements. These are subsequently formatted as "theorem name: informal statement" and incorporated into an informal corpus. Figure 3 shows the prompt used for informalizing the theorem Exists.choose_spec, including the formal statement, documentation string, and the definition of Exists.choose as annotations to aid the LLM’s understanding.
Notably, the authors of [7] also employed an LLM for informalizing mathlib4 statements, relying solely on the formal statements. We argue that providing the LLM with additional context, such as related definitions and documentation strings, enhances its ability to accurately interpret and convert formal theorems into their informal counterparts.
3.2 Semantic Search Engine for Mathlib4
After obtaining the informal corpus, we employ dense embedding models, which excel at capturing semantic information, to encode the informal-formal theorem pairs. Recent advancements in text embedding models have introduced the practice of integrating specific task instructions into either queries or documents, enhancing the model’s adaptability to diverse tasks and boosting performance in zero-shot settings [28, 13, 31]. Consequently, for the purpose of theorem retrieval within mathlib4, we enrich our corpus documents with specific instructions. We adopt the following instruction template:
[…]
Here, {document} denotes
[…]
These task-specific instructions significantly influence the performance of embedding models. We will investigate the effect of varying task instructions on the overall retrieval performance in Section 5.3.1.
Embedding the entire corpus, although a time-consuming process, is performed offline and does not require repetition for each use of the engine. We utilize Chroma DB to store the embeddings. Upon receiving a query, our system vectorizes it and retrieves theorems based on their cosine distance to the query in the embedding space. Chroma DB employs the Hierarchical Navigable Small World (HNSW) algorithm [15], an efficient approximate nearest neighbor search method, ensuring rapid retrieval from the corpus.
3.3 Query Augmentation
To enhance query clarity and achieve a more accurate embedding, our system incorporates a query augmentation step. As illustrated in Figure 4, this process involves prompting an LLM101010We use gpt-4-0125-preview for query augmentation. to transform a brief, vague query into a detailed statement that includes both informal and formal statements, ensuring mathematical equivalence with the original query. We guide the LLM with specific principles for query augmentation, emphasizing the importance of precision, the use of LaTeX for mathematical expressions, and the clarification of ambiguous inputs. Additionally, we provide examples of query augmentations to improve the LLM’s comprehension of the task. Although Lean 4 code generation by LLMs may occasionally introduce inaccuracies due to the limited presence of mathlib4 in their training data, this approach effectively enriches the query with additional contextual information.
[…]
This formatted query is then vectorized and utilized by the search engine to retrieve theorems based on the cosine distance in the embedding space.
4 Mathlib4 Semantic Search Benchmark
To rigorously assess and compare the efficacy of various retrieval methods, we have established the Mathlib4 semantic search benchmark. This benchmark encompasses a curated set of queries, relevance labels of each each Mathlib4 theorem for these queries, and a collection of performance metrics. In this section, we will explain this benchmark in detail. Section 4.1 describes the composition of the query set. In Section 4.2, we explain the relevance criteria for our benchmark along with the whole labeling procedure. Section 4.3 lists the performance metrics used in our benchmark.
4.1 Composition of the Query set
The query set of our benchmark has 50 distinct queries, spanning various mathematical disciplines including calculus, abstract algebra, linear algebra, number theory, algebraic number theory, set theory, and mathematical logic. This selection aims to cover a broad spectrum of topics and complexities. We argue that this is a common size for mathematical information retrieval database, as evidenced by the ARQMath1, ARQMath2, and ARQMath3 databases from the MIR field, which contain 77, 71, and 78 queries, respectively, for their answer retrieval tasks [16].
To optimize labeling efforts, we organize queries with identical search intents into 18 distinct, non-overlapping groups, each containing at least two distinct queries. This approach assumes that all queries within a group have the same relevance score for any document and significantly reduces the need for repetitive labeling, as each document is evaluated just once per query group. Furthermore, to provide a more detailed assessment and mitigate the impact of duplicate document labels, we consider four prevalent forms of mathematical queries: natural language descriptions, LaTeX formulas, theorem names, and Lean 4 term descriptions. In each query group, we strive to include as many different description forms as possible. Table 1 provides statistics and examples of the query set. We note that not all query groups have all four different representation forms, as it is often the case that a theorem may not have an official name or its representation in a LaTeX formula might be redundant.
| Category | Count | Example 1 | Example 2 |
Natural Description |
18 |
If there exist injective maps of sets from to and from to , then there exists a bijective map between and . |
If implies , then not implies not . |
LaTeX Formula |
15 |
If there exist injective, injective, then there exists bijective. |
Theorem Name |
7 |
Schroeder Bernstein Theorem |
Modus Tollens |
Lean 4 Term |
10 |
{f : A → B} {g : B → A} (hf : Injective f) (hg : Injective g) : h, Bijective h | (p → q) → (¬q → ¬p) |
4.2 Relevance Judgments
During the process of collecting relevance labels for query-document pairs, we adopt an approach similar to ARQMath[16]. Initially, we establish a carefully crafted relevance assessment criteria, elaborated in Table 2. While doing labeling, assessors are instructed to evaluate whether a given search result facilitates their mathematical formalization workflow. Instead of simply considering similarity in topic, presented form, formula structure or mathematical induction relationship, the relevance we consider here are deeply engaged with Lean 4 experts’ need.
| Rating | Label | Score | Definition |
Exact Match |
2 |
1 |
Exact match to the query or being a stronger statement |
Relevant |
1 |
0.3 |
Useful in locating where the corresponding statement should be |
Irrelevant |
0 |
0 |
Not expected to be useful in mathematics formalization workflow |
For each set of queries grouped by identical search intentions, assessors are presented with the top 50 theorems list retrieved by an intermediate version of our search engine. In addition to evaluating the provided theorems, assessors are tasked with identifying and adding any relevant theorems that may have been omitted from the initial list by inspecting the files where "Exact Match" items in the list are located, with particular attention to those with "Exact Match" rating.
Given the structured organization of Mathlib4, where related theorems are often located within the same file, it is reasonable to assume that any items not in the list are irrelevant to this query group. We also note that all queries are assured to have at least one exact match in Mathlib4. These two assumptions support the performance metrics we use in the benchmark, as described in the following subsection.
4.3 Performance Measures
To compare different retrieval paradigms’ performance on our labeled dataset, three commonly used metrics are adopted in our benchmark. Precision@k calculates the average relevant document proportion in top k retrieved results among the query set :
where stands for the retrieved theorem list for i-th query, and if and only if the j-th retrieved result of i-th query is "Exact Match", otherwise . Based on the assumption that all unlabeled theorems are irrelevant, it is also reasonable to calculate the Recall@k:
where is the number of "Exact Match" theorems for i-th query.
The third metric, nDCG (normalized Discounted Cumulative Gain), incorporates the position of the retrieved result. A decaying weight is allocated to each position:
where stands for the score of j-th retrieved theorem with respect to i-th query, defined in Table 2. We note that our score is not proportional to label number, aiming to emphasize the importance of exact matching in mathlib4 retrieval task. We further define as the highest possible by properly arranging retrieved theorems, and finally:
To sum up, nDCG@k measure both the retrieving and ranking ability of the given engine in a unified and detailed way. Meanwhile, P@10 (Precision@10) and R@10 (Recall@10) focus on the retrieval effectiveness solely. In our benchmark, all three metrics will be reported both on the whole query set and on each query category.
5 Experiments
In this section, we evaluate the performance of various theorem retrieval methods using our benchmark. The experimental configurations, including embedding models, task instructions and computational expenses, are detailed in Section 5.1. We compare the performances of different retrieval methodologies in Section 5.2. Section 5.3 analyzes the impact of different task instructions, the documents content type, and the query augmentations employed in our approach.
5.1 Experiment setup
We have considered four embedding models in our experiments: text-embedding-ada-002 and text-embedding-3-large from OpenAI, UAE-Large-V1[13], and [31]. We use default task instructions in UAE-Large-V1. For , asymmetrical task instructions were employed in the non-augmented query setting, and symmetrical task instructions were applied in the query augmentation setting, as shown in Table 3. We observe that the original implementation of does not use document-side task instructions to reduce computational cost during document indexing in multiple retrieval tasks. Considering the complexity of the mathematical embedding task, we have modified this to a two-sided prompt approach, as has shown substantial performance degradation with one-sided prompt setting in our benchmarks.
| Query aug. & Doc Type | Side | Input with Task Instructions |
None & Formal |
Query |
[…]
Doc |
[…]
F+IF & F+IF |
Query |
[…]
Doc |
[…]
For the baseline models, Moogle, along with all four embedding models applied on original Lean 4 formal corpus and unaugmented query, are used in our experiments. The same four embedding models equipped with formal + informal query augmentation and formal + informal document corpus are also tested on our benchmark to demonstrate the efficacy of our approach. During evaluation, inputs to the embedding model exceeding 4096 characters were truncated due to GPU memory limitations. For all tested models, the corpus embedding process was completed within three hours on a single V100 GPU. Subsequently, evaluations of these models were conducted using the three metrics mentioned earlier.
5.2 Main Results
| Model | Corpus | Query aug. | nDCG@20 | P@10 | R@10 |
| Baselines | |||||
| Moogle† | \ | \ | 0.365† | 0.092† | 0.513† |
| OpenAI v2 | F | \ | 0.312 | 0.078 | 0.405 |
| OpenAI v3 | F | \ | 0.493 | 0.128 | 0.622 |
| UAE-Large-V1 | F | \ | 0.233 | 0.066 | 0.307 |
| F | \ | 0.593 | 0.132 | 0.687 | |
| Our Methods | |||||
| OpenAI v2 | F+IF | F+IF | 0.553 | 0.144 | 0.707 |
| OpenAI v3 | F+IF | F+IF | 0.691 | 0.178 | 0.837 |
| UAE-Large-V1 | F+IF | F+IF | 0.368 | 0.084 | 0.440 |
| F+IF | F+IF | 0.733 | 0.196 | 0.913 | |
| Model | nDCG@20 | P@10 | ||||||
| ND | LF | TN | LT | ND | LF | TN | LT | |
| Baselines | ||||||||
| Moogle† | 0.369† | 0.324† | 0.333† | 0.441† | 0.083† | 0.107† | 0.071† | 0.100† |
| OpenAI v2 | 0.276 | 0.379 | 0.000 | 0.498 | 0.061 | 0.107 | 0.000 | 0.120 |
| OpenAI v3 | 0.479 | 0.553 | 0.235 | 0.610 | 0.122 | 0.160 | 0.057 | 0.140 |
| UAE-Large-V1 | 0.301 | 0.216 | 0.004 | 0.298 | 0.078 | 0.067 | 0.000 | 0.090 |
| 0.576 | 0.633 | 0.294 | 0.774 | 0.139 | 0.140 | 0.043 | 0.170 | |
| Our Methods | ||||||||
| OpenAI v2 | 0.536 | 0.533 | 0.571 | 0.600 | 0.111 | 0.160 | 0.186 | 0.150 |
| OpenAI v3 | 0.681 | 0.657 | 0.772 | 0.703 | 0.167 | 0.180 | 0.200 | 0.180 |
| UAE-Large-V1 | 0.415 | 0.337 | 0.371 | 0.329 | 0.100 | 0.080 | 0.071 | 0.070 |
| 0.748 | 0.712 | 0.855 | 0.654 | 0.194 | 0.200 | 0.214 | 0.180 | |
Table 4 and 5 present the results of Moogle and four embedding models. , when integrated with our retrieval pipeline, achieves the best performance across three overall metrics and significantly outperforms all other methods, including its own performance on a formal corpus with unaugmented queries. We observe that all retrieval methods benefit from using an augmented corpus and queries, as augmentation expands concrete mathematics and Lean 4 terms, and these embedding models typically perform better on symmetrical retrieval tasks [31].
Examining the results averaged by category, we find that attains the best results in nearly all categories, with the exception of nDCG@20 for Lean 4 Terms. In this case, the augmented informal information may impair retrieval performance, as it is not essential in formal-formal retrieval, which focuses on lexical matches, but it does alter the final embedding representation. Meanwhile, our method markedly enhances performance in the Theorem Name category, because the expansion of the statement for a given theorem name is crucial for successful retrieval in this category. Our approach of informalization and query augmentation facilitates this process for embedding models.
5.3 Ablation Studies
We demonstrate that our pipeline enables embedding models to unpack concrete mathematical concepts, transforming the retrieval task into a symmetric one and thereby enhancing their performance in Mathlib4 retrieval tasks. In this subsection, we analyze our results and conduct ablation studies from various perspectives. The necessity of proper task instructions is illustrated in Section 5.3.1, while the impact of document preparations and query augmentation is examined in Sections 5.3.2 and 5.3.3, respectively.
5.3.1 Necessity of proper task instructions
Recent models incorporate task instructions into their training to enhance zero-shot performance [28, 31]. In this part, we analyze the impact of task instructions in mathematical informal retrieval. We present a visualization of the embeddings of all 50 queries, generated by four embedding engine in our benchmark, using t-SNE[29]. We use default task instruction for UAE-Large-V1. For , in addition to using empty instruction, we also test mathematics retrieving instruction "Given a math search query, retrieve theorems mathematically equivalent to the query" and Lean 4 retrieving instruction "Given a math search query, retrieve Lean 4 written theorems mathematically match the query".
As illustrated in Figure 5, the cluster of theorem names appear in the results produced by all four embedding models. However, these clusters vanish when we provide with mathematics-aware task instructions. Given that we assume all queries within the same query group (denoted by identical colors in Figure 5) as having the same search intent, they should be proximal on the graph, as t-SNE maintains the relative distance relationships between vectors. Thus, the more effective the embedding engine, the more closely the dots of the same color group together. A comparison of subfigures 4(c) and 4(d) with 4(a) and 4(b) reveals that text-embedding-3-large and demonstrate superior performance in discerning mathematical search intentions, aligning with our main results (Table 4).
The presence of theorem name clusters in these subfigures, however, contradicts this principle, suggesting that the embedding models perceive these theorem names as more similar to each other rather than correctly associating them with their respective query groups. While this might be acceptable for other retrieval tasks, it is clearly inappropriate for our mathematical information retrieval context. In contrast, the absence of theorem clusters in subfigures 4(e) and 4(f) suggests that these embedding methods successfully group theorem names with their corresponding statements, despite their notable differences in appearance and structure. This indicates their proficiency in comprehending mathematical theorems and recognizing search intents. In summary, appropriate task instructions enhance the embedding models’ sensitivity to the search intents of mathematical concepts, thereby improving retrieval effectiveness on Mathlib4.
5.3.2 Ablation of document preparations
| Model & Corpus | nDCG@20 | P@10 | R@10 | ||||
| ND | LF | TN | LT | All | All | All | |
| without query augmentation | |||||||
| OpenAI v3 large | |||||||
| formal corpus | 0.479 | 0.553 | 0.235 | 0.610 | 0.493 | 0.128 | 0.622 |
| informal corpus | 0.591 | 0.527 | 0.365 | 0.451 | 0.512 | 0.134 | 0.628 |
| formal+informal corpus | 0.607 | 0.549 | 0.267 | 0.622 | 0.545 | 0.140 | 0.677 |
| formal corpus | 0.576 | 0.633 | 0.294 | 0.774 | 0.593 | 0.132 | 0.687 |
| informal corpus | 0.696 | 0.636 | 0.616 | 0.604 | 0.648 | 0.174 | 0.773 |
| formal+informal corpus | 0.749 | 0.687 | 0.698 | 0.612 | 0.701 | 0.184 | 0.845 |
| with query augmentation | |||||||
| OpenAI v3 large | |||||||
| formal corpus | 0.637 | 0.617 | 0.565 | 0.647 | 0.623 | 0.160 | 0.747 |
| informal corpus | 0.603 | 0.650 | 0.810 | 0.575 | 0.640 | 0.166 | 0.783 |
| formal+informal corpus | 0.681 | 0.657 | 0.772 | 0.703 | 0.691 | 0.178 | 0.837 |
| formal corpus | 0.688 | 0.685 | 0.631 | 0.694 | 0.680 | 0.178 | 0.847 |
| informal corpus | 0.711 | 0.646 | 0.824 | 0.670 | 0.699 | 0.184 | 0.877 |
| formal+informal corpus | 0.748 | 0.712 | 0.855 | 0.654 | 0.733 | 0.196 | 0.913 |
| Query aug. & Doc Type | Side | Task Instructions |
None & Informal |
Query |
Given a math search query, retrieve theorems mathematically equivalent to the query |
Doc |
Represent the given math theorem statement for retrieving related statement by natural language query | |
None & F+IF |
Query |
Given a math search query, retrieve theorems stated in bilingual Lean 4 + natural language that mathematically match the query |
Doc |
Represent the given formal math statement written in Lean 4 concatenated with its natural language explanation for retrieving related statement by natural language query | |
Formal & Formal |
Query & Doc |
Retrieve math theorems stated in Lean 4 that are mathematically equivalent to the given one |
IF & IF |
Query & Doc |
Retrieve math theorems that are mathematically equivalent to the given one |
Table 6 presents the effects of changing corpus type with or without query augmentation. We simply use the original Lean 4 statement and the informalized statement in our formal + informal corpus to form the formal and informal corpus. Under no query augmentation setting, results indicate that both engines underperform with incomplete corpus components, yielding lower overall scores on all metrics. Specifically, the formal + informal corpus enhances performance on Lean 4 Terms by incorporating formal data and improves score on the Natural Description category by providing hybrid information. For the Theorem Name category, significantly benefits from the formal + informal corpus by utilizing the unfolded mathematical descriptions provided by the informalized statement, whereas text-embedding-3-large shows diminished results, likely due to the lack of training on mixed-domain texts and absence of adaptive instructions. We note that formal corpus outperforms formal+informal corpus on Lean 4 Terms category for , with the same reason we explained in Section 5.2: retrieval on Lean 4 Terms highly relies on lexical instead of semantic information, and the existence of informal information alters the final embedding vector. Similar results can be observed when using augmented queries.
When using non-augmented queries, a more detailed examination of the results that takes averages on query groups, further validates our analysis. As depicted in Figure 6, in the majority of instances, the "formal + informal" configuration does not experience the abrupt decline in performance observed in solely formal or informal settings. The informalized statement provided in the corpus mitigate the lexical mismatch between query and formal statement. Furthermore, the incorporation of both informal and formal information occasionally leads to enhanced outcomes. These two arguments support the choice of hybrid corpus in our pipeline.
5.3.3 Ablation of query augmentations
The effect of query augmentation can be observed by comparing the results in the upper and lower sections of Table 6. Query augmentation is designed to align the types of queries and documents; thus, we use the formal and informal part of our augmented queries for the formal and informal corpora, respectively. As indicated in the table, augmented queries enhance the performance of all methods across the board. Specifically, query augmentation markedly improves the performance of all methods in identifying Theorem Names by expanding terms to provide a richer context. We claim that the process of informalization serves a similar purpose, and that such additional information can be extracted through the use of appropriate instructions. This is evidenced by the high nDCG@20 scores achieved by on Theorem Names by using informal or formal + informal corpus, under the setting of using non-augmented query. The comparable overall performance of on formal + informal corpus in this case, presents it as a cost-effective option for frequent usage due to its potential to reduce LLM API costs.
6 Conclusion
In this paper, we introduce a semantic search engine designed to enable users to locate theorems in mathlib4 using informal queries. Specifically, we translate the formal statements of mathlib4 theorems into informal versions and develop our search engine to work with a corpus of informal-formal theorem pairs. Additionally, we construct a dataset to facilitate evaluation. Our comprehensive experiments on this dataset reveal that the best theorem retrieval performance is attained by augmenting the user’s query appropriately and embedding the content of the corpus simultaneously. Consequently, our system employs a strategy that first augments the query, followed by a semantic search, thereby precisely aligning with the users’ search intentions.
Our future research will focus on three primary directions. First, in terms of informalizing mathlib4, we aim to design guidelines for translation and provide examples of converting mathlib4 statements to informal language for LLMs, enhancing the informal corpus’s quality. Second, regarding the mathlib4 semantic search benchmark, we plan to continually enlarge the query set, aiming for a more comprehensive benchmark. Lastly, for the semantic search engine itself, we intend to fine-tune a text embedding model on the task of theorem retrieval, aiming to improve the search engine’s performance.